# Distributionally Robust Receive Beamforming

### v2

@@ IV-B1 Wasserstein Distributionally Robust Beamforming
The first constraint in the above display is a particularization of the Wasserstein distance between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$. Problem (46) is equivalent, in the sense of the same maximizer(s), to
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},\end{array}$ (47)
since the objective of (46) is increasing in $\bm{R}$; cf. Theorem 1.
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{U}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2\bm{U}}\right]\leq\epsilon^{2}_{0}\\ &\left[\begin{array}[]{cc}\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2}&\bm{U}\\ \bm{U}&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (48)
@@ IV-B2 F-Norm Distributionally Robust Beamforming
Under the $F$-norm, we just need to replace the Wasserstein distance. To be specific, for example, (47) becomes
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}(\bm{R}-\hat{\bm{R}})}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0}.\end{array}$ (52)
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{U}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{U}}\right]\leq\epsilon^{2}_{0},\\ &\left[\begin{array}[]{cc}\bm{U}&(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}\\ (\bm{R}-\hat{\bm{R}})&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0},\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (53)
@@ IV-C Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener-MF}}&=[\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{R}_{x}+\lambda\bm{I}_{N}]^{-1}\\ &=[\bm{R}_{s}\bm{H}^{\mathsf{H}}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\lambda\bm{I}_{N}]^{-1},\end{array}$ (55)
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Wiener-MF}}=[\hat{\bm{R}}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}]\cdot[\hat{\bm{R}}_{x}+\lambda\bm{I}_{N}+\epsilon_{0}\bm{I}_{N}]^{-1},\end{array}$
@@ V Distributionally Robust Nonlinear Estimation
For the convenience of the technical treatment, we study the estimation problem in real spaces. Nonlinear estimators are to be limited in reproducing kernel Hilbert spaces and feedforward multi-layer neural network function spaces.
@@ V-A1 General Framework and Concrete Examples
where $\hat{\bm{R}}_{\underline{z}}=\frac{1}{L}\bm{K}^{2}$, $\hat{\bm{R}}_{\underline{zs}}=\frac{1}{L}\bm{K}\underline{\bm{S}}^{\mathsf{T}}$, $\hat{\bm{R}}_{\underline{s}}=\frac{1}{L}\underline{\bm{S}}\underline{\bm{S}}^{\mathsf{T}}$, and $\underline{\bm{S}}\coloneqq[\operatorname{Re}\bm{S};~{}\operatorname{Im}\bm{S}]=[\underline{\bm{s}}_{1},\underline{\bm{s}}_{2},\ldots,\underline{\bm{s}}_{L}]$. In addition, the strong min-max property holds for (58): i.e., the order of $\min$ and $\max$ can be exchanged provided that the first constraint is compact convex. As a result, given every pair of $(\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}})$, the optimal Wiener beamformer is [^3] For numerical stability in the inversion of $\bm{K}$, $L$ should not be too large.
$\bm{W}^{\star}_{\text{RKHS}}=\hat{\bm{R}}^{\mathsf{T}}_{\underline{zs}}\cdot\hat{\bm{R}}^{-1}_{\underline{z}}=\underline{\bm{S}}\bm{K}\bm{K}^{-2}=\underline{\bm{S}}\bm{K}^{-1},$ (59)
@@
Treating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ as, or approximating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ using, a joint Gaussian random vector due to the linear estimation relation $\hat{\underline{\mathbf{s}}}=\bm{W}\underline{\mathbf{z}}$ in RKHS [cf. (57)], then the results in Lemma 1 apply. For details, see Appendix F of the online supplementary materials. $\square$
@@ V-A2 Multi-Frame Case: Dynamic Channel Evolution
The solution of (67) is given by (cf. (59))
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{RKHS-MF}}&=[\bm{R}_{\underline{zs}}+\lambda\bm{W}^{\prime\mathsf{T}}][\bm{R}_{\underline{z}}+\lambda\bm{I}_{L}]^{-1}\\ &=\left(\frac{1}{L}\underline{\bm{S}}\bm{K}+\lambda\bm{W}^{\prime\mathsf{T}}\right)\cdot\left(\frac{1}{L}\bm{K}^{2}+\lambda\bm{I}_{L}\right)^{-1},\end{array}$ (68)
@@ V-B Neural Networks
where $\bm{W}_{[R]}\coloneqq\{\bm{W}_{1},\bm{W}_{2},\ldots,\bm{W}_{R}\}$ and $\bm{\phi}_{\bm{W}_{[R]}}(\underline{\mathbf{x}})$ is defined in (4). Problem (69) is highly nonlinear in both argument $\underline{\mathbf{x}}$ and parameter $\bm{W}_{[R]}$, which is different from the case in reproducing kernel Hilbert spaces where the $\bm{W}_{[R]}$-linearization features. Hence, problem (69) is too complicated to solve to global optimality. According to [27, Cor. 33], under several technical conditions (plus the boundedness of the feasible region of $\bm{W}_{[R]}$), (69) is upper bounded by a spectral-norm-regularized empirical risk minimization problem
@@
for some regularization coefficient $\lambda^{\prime}\geq 0$, where $\|\cdot\|_{2}$ denotes the spectral norm of a matrix (i.e., the induced $2$-norm). Eq. (70) rigorously justifies the popular norm regularization method in training neural networks: By reducing the upper bound (70) of (69), the value of (69) can be diminished as well. The regularized ERM problem (70) is reminiscent of the ridge regression and the kernel ridge regression methods in Theorems 2 and 3 for distributional robustness in linear regression and RKHS linear regression, respectively. Supposing that $\bm{W}^{\star}_{[R]}$ is an (approximated, or sub-optimal) solution [^4] Neural networks are hard to be globally optimized. of (70), then the distributionally robust optimal estimator of the transmitted signal $\mathbf{s}$ can be obtained as
@@ VI Experiments
We consider a point-to-point multiple-input-multiple-output (MIMO) wireless communication problem where the transmitter is located at $[0,0]$ and the receiver is at $[500\text{m},450\text{m}]$. We randomly sample $25$ points according to the uniform distribution on the square of $[0,500\text{m}]\times[0,500\text{m}]$ to denote the scatters’ positions; i.e., there exist $25$ radio paths. All the source data and codes are available online at GitHub with thorough implementation comments: https://github.com/Spratm-Asleaf/Beamforming. In this section, we only present major experimental setups and results; readers can use the shared source codes to explore (or verify) minor ones.
The following eleven methods are implemented in the experiments: 1) Wiener: Wiener beamformer (13), upper expression; 2) Wiener-DL: Wiener beamformer with diagonal loading (30), upper expression; 3) Wiener-DR: Distributionally robust Wiener beamformer (49) and (53); 4) Wiener-CE: Channel-estimation-based Wiener beamformer (13), lower expression; 5) Wiener-CE-DL: Channel-estimation-based Wiener beamformer with diagonal loading (30), lower expression; 6) Wiener-CE-DR: Distributionally robust channel-estimation-based Wiener beamformer (41) and (31); 7) Capon: Capon beamformer (38) for $\epsilon_{0}=0$; 8) Capon-DL: Capon beamformer with diagonal loading (38); 9) ZF: Zero-forcing beamformer where $\bm{W}_{\text{ZF}}\coloneqq(\hat{\bm{H}}^{\mathsf{H}}\hat{\bm{H}})^{-1}\hat{\bm{H}}^{\mathsf{H}}$ and $\hat{\bm{H}}$ denotes the estimated channel matrix; 10) Kernel: Kernel receiver (61) with $\epsilon_{0}=0$ in (60); and 11) Kernel-DL: Kernel receiver with diagonal loading (65). Note that the diagonal-loading-based methods are particular cases of distributionally robust beamformers; see, e.g., Corollary 1 and Example 4. The deep-learning-based (DL-based) methods in Subsection V-B are not implemented in this section because they have been deeply studied in our previous publications, e.g., [10, 12]; we only comment on the advantages and disadvantages of DL-based methods compared with the listed eleven methods in Section VII (Conclusions).
@@
The beamformers are determined on the training data set (i.e., pilot data). The performance evaluation method of beamformers is mean-squared estimation error (MSE) on the test data set (i.e., non-pilot communication data): to be specific, $\|\bm{S}_{\text{test}}-\hat{\bm{S}}_{\text{test}}\|^{2}_{F}/(M\times L_{\text{test}})$ where $\bm{S}_{\text{test}}\in\mathbb{C}^{M\times L_{\text{test}}}$ is the test data block, $\hat{\bm{S}}_{\text{test}}$ is its estimate, and $L_{\text{test}}$ is the length of non-pilot test data units. As data-driven machine learning methods, all parameters (e.g., uncertainty quantification coefficients $\epsilon$’s) of beamformers can be tuned using the popular cross-validation (e.g., one-shot cross-validation) method. The parameters can also be empirically tuned to save training times because cross-validation imposes a significant computational burden. This article mainly uses the empirical tuning method (i.e., trial-and-error) to tune each beamformer to achieve its best average performance. For each test case, the MSE performances are averaged on $250$ Monte–Carlo episodes.
@@
    Since the signal model under impulse channel noises is no longer linear Gaussian, the optimal beamformer in the MSE sense must be nonlinear. Therefore, the Kernel and the Kernel-DL methods have the potential to outperform other linear beamformers, i.e., to suppress outliers. However, in practice, the non-robust Kernel method may undergo numerical instability in calculating the inverse of the kernel matrix $\bm{K}$. Therefore, its actual MSEs are not necessarily smaller than those of linear beamformers. Nevertheless, the robust Kernel-DL method consistently outperforms all other beamformers.
@@
    Distributionally robust beamformers (including diagonal-loading ones) can combat the adverse effect introduced by the limited pilot size and several types of uncertainties in the signal model (e.g., outliers). To be specific, for example, all diagonal-loading beamformers can outperform their original non-diagonal-loading counterparts; cf. the Wiener and the Wiener-DL methods, the Wiener-CE and the Wiener-CE-DL methods, the Capon and the Capon-DL methods, and the Kernel and the Kernel-DL methods. In addition, the Wiener-DR beamformer (53) using the $F$-norm uncertainty set has the potential to outperform the Wiener-DL beamformer (30) that employs the simple uncertainty set (26).
@@
Remarks on Parameter Tuning: From experiments, we find that the uncertainty quantification coefficients $\epsilon$’s (e.g., in diagonal loading) can be neither too large nor too small. When $\epsilon$’s are too large, the beamformers become overly conservative, while when $\epsilon$’s are too small, the beamformers cannot offer sufficient robustness against data scarcity and model uncertainties. In both cases of inappropriate $\epsilon$’s, the performances of beamformers degrade significantly. Therefore, $\epsilon$’s must be carefully tuned in practice, and a rigorous method to tune $\epsilon$’s can be the cross-validation method on the training data set (i.e., the pilot data set). If practitioners just pursue satisfaction rather than optimality, the empirical tuning method is recommended to save training times.
@@
Beamformer MSE Time Beamformer MSE Time
@@
Beamformer MSE Time Beamformer MSE Time
@@
Beamformer MSE Time Beamformer MSE Time
@@
Beamformer MSE Time Beamformer MSE Time
@@
Beamformer MSE Time Beamformer MSE Time
@@
Beamformer MSE Time Beamformer MSE Time

### v3

@@ IV-B1 Wasserstein Distributionally Robust Beamforming
The first constraint in the above display is a particularization of the Wasserstein distance between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$.
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{V},\bm{U}}&\operatorname{Tr}[\bm{R}_{s}-\bm{V}]\\ \text{s.t.}&\left[\begin{array}[]{cc}\bm{V}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x}\end{array}\right]\succeq\bm{0}\\ &\operatorname{Tr}\left[{\bm{R}+\hat{\bm{R}}-2\bm{U}}\right]\leq\epsilon^{2}_{0}\\ &\left[\begin{array}[]{cc}\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2}&\bm{U}\\ \bm{U}&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{V}\succeq\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (48)
@@ IV-B2 F-Norm Distributionally Robust Beamforming
Under the $F$-norm, we just need to replace the Wasserstein distance. To be specific, (47) becomes
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}(\bm{R}-\hat{\bm{R}})}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0}.\end{array}$ (52)
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{V},\bm{U}}&\operatorname{Tr}[\bm{R}_{s}-\bm{V}]\\ \text{s.t.}&\left[\begin{array}[]{cc}\bm{V}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x}\end{array}\right]\succeq\bm{0}\\ &\operatorname{Tr}\left[{\bm{U}}\right]\leq\epsilon^{2}_{0},\\ &\left[\begin{array}[]{cc}\bm{U}&(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}\\ (\bm{R}-\hat{\bm{R}})&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0},\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{V}\succeq\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (53)
@@ IV-C Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener-MF}}&=[\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}]^{\mathsf{H}}[\bm{R}_{x}+\lambda\bm{I}_{N}]^{-1}\\ &=[\bm{R}_{s}\bm{H}^{\mathsf{H}}+\lambda\bm{W}^{\prime}][\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\lambda\bm{I}_{N}]^{-1},\end{array}$ (55)
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Wiener-MF}}=[\hat{\bm{R}}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}]^{\mathsf{H}}\cdot[\hat{\bm{R}}_{x}+\lambda\bm{I}_{N}+\epsilon_{0}\bm{I}_{N}]^{-1},\end{array}$
@@ V Distributionally Robust Nonlinear Estimation
For the convenience of the technical treatment, we study the estimation problem in real spaces. Nonlinear estimators, which are suitable for non-Gaussian $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, are to be limited in reproducing kernel Hilbert spaces and feedforward multi-layer neural network function spaces.
@@ V-A1 General Framework and Concrete Examples
where $\hat{\bm{R}}_{\underline{z}}=\frac{1}{L}\bm{K}^{2}$, $\hat{\bm{R}}_{\underline{zs}}=\frac{1}{L}\bm{K}\underline{\bm{S}}^{\mathsf{T}}$, $\hat{\bm{R}}_{\underline{s}}=\frac{1}{L}\underline{\bm{S}}\underline{\bm{S}}^{\mathsf{T}}$, and $\underline{\bm{S}}\coloneqq[\operatorname{Re}\bm{S};~{}\operatorname{Im}\bm{S}]=[\underline{\bm{s}}_{1},\underline{\bm{s}}_{2},\ldots,\underline{\bm{s}}_{L}]$. In addition, the strong min-max property holds for (58): i.e., the order of $\min$ and $\max$ can be exchanged provided that the first constraint is compact convex. As a result, given every pair of $(\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}})$, the optimal Wiener beamformer is
$\bm{W}^{\star}_{\text{RKHS}}=\bm{R}^{\mathsf{T}}_{\underline{zs}}\cdot\bm{R}^{-1}_{\underline{z}}$ (59)
@@
Treating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ as, or approximating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ using, a joint Gaussian random vector due to the linear estimation relation $\hat{\underline{\mathbf{s}}}=\bm{W}\underline{\mathbf{z}}$ in RKHS [cf. (57)], then the results in Lemma 1 apply. For details, see Appendix G. $\square$
@@ V-A2 Multi-Frame Case: Dynamic Channel Evolution
The solution to (67) is given by (cf. (59))
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{RKHS-MF}}&=[\bm{R}_{\underline{zs}}+\lambda\bm{W}^{\prime\mathsf{T}}]^{\mathsf{T}}[\bm{R}_{\underline{z}}+\lambda\bm{I}_{L}]^{-1}\\ &=\left(\frac{1}{L}\underline{\bm{S}}\bm{K}+\lambda\bm{W}^{\prime}\right)\cdot\left(\frac{1}{L}\bm{K}^{2}+\lambda\bm{I}_{L}\right)^{-1},\end{array}$ (68)
@@ V-B Neural Networks
where $\bm{W}_{[R]}\coloneqq\{\bm{W}_{1},\bm{W}_{2},\ldots,\bm{W}_{R}\}$ and $\bm{\phi}_{\bm{W}_{[R]}}(\underline{\mathbf{x}})$ is defined in (4). Problem (69) is highly nonlinear in both argument $\underline{\mathbf{x}}$ and parameter $\bm{W}_{[R]}$, which is different from the case in reproducing kernel Hilbert spaces where the $\bm{W}$-linearization features. Hence, problem (69) is too complicated to solve to global optimality. According to [27, Cor. 33], under several technical conditions (plus the boundedness of the feasible region of $\bm{W}_{[R]}$), (69) is upper bounded by a spectral-norm-regularized empirical risk minimization problem
@@
for some regularization coefficient $\lambda^{\prime}\geq 0$, where $\|\cdot\|_{2}$ denotes the spectral norm of a matrix (i.e., the induced $2$-norm). Eq. (70) rigorously justifies the popular norm regularization method in training neural networks: By diminishing the upper bound (70) of (69), the true error in (69) can be controlled from above. The regularized ERM problem (70) is reminiscent of the ridge regression and the kernel ridge regression methods in Theorems 2 and 3 for distributional robustness in linear regression and RKHS linear regression, respectively. Supposing that $\bm{W}^{\star}_{[R]}$ is an (approximated, or sub-optimal) solution [^3] Neural networks are hard to be globally optimized. of (70), then the distributionally robust optimal estimator of the transmitted signal $\mathbf{s}$ can be obtained as
@@ VI Experiments
We consider a point-to-point multiple-input-multiple-output (MIMO) wireless communication problem where the transmitter is located at $[0,0]$ and the receiver is at $[500\text{m},450\text{m}]$. We randomly sample $25$ points according to the uniform distribution on the square of $[0,500\text{m}]\times[0,500\text{m}]$ to denote the scatters’ positions; i.e., there exist $26$ radio paths. All the source data and codes are available online at GitHub with thorough implementation comments: https://github.com/Spratm-Asleaf/DRRC. In this section, we only present major experimental setups and results; readers can use the shared source codes to explore (or verify) minor ones.
The following eleven methods are implemented in the experiments: 1) Wiener: Wiener beamformer (12), upper expression; 2) Wiener-DL: Wiener beamformer with diagonal loading (30), upper expression; 3) Wiener-DR: Distributionally robust Wiener beamformer (49) and (53); 4) Wiener-CE: Channel-estimation-based Wiener beamformer (12), lower expression; 5) Wiener-CE-DL: Channel-estimation-based Wiener beamformer with diagonal loading (30), lower expression; 6) Wiener-CE-DR: Distributionally robust channel-estimation-based Wiener beamformer (42) and (31); 7) Capon: Capon beamformer (39) for $\epsilon_{0}=0$; 8) Capon-DL: Capon beamformer with diagonal loading (39); 9) ZF: Zero-forcing beamformer where $\bm{W}_{\text{ZF}}\coloneqq(\hat{\bm{H}}^{\mathsf{H}}\hat{\bm{H}})^{-1}\hat{\bm{H}}^{\mathsf{H}}$ and $\hat{\bm{H}}$ denotes the estimated channel matrix; 10) Kernel: Kernel receiver (61) with $\epsilon_{0}=0$ in (60); and 11) Kernel-DL: Kernel receiver with diagonal loading (65). Note that the diagonal-loading-based methods are particular cases of distributionally robust combiners; see, e.g., Corollary 1 and Example 4. The deep-learning-based (DL-based) methods in Subsection V-B are not implemented in this section because they have been deeply studied in our previous publications, e.g., [10, 12]; we only comment on the advantages and disadvantages of DL-based methods compared with the listed eleven methods in Section VII (Conclusions).
@@
The combiners are determined on the training data set (i.e., pilot data). The performance evaluation method of combiners is mean-squared estimation error (MSE) on the test data set (i.e., non-pilot communication data): to be specific, $\|\bm{S}_{\text{test}}-\hat{\bm{S}}_{\text{test}}\|^{2}_{F}/(M\times L_{\text{test}})$ where $\bm{S}_{\text{test}}\in\mathbb{C}^{M\times L_{\text{test}}}$ is the test data block, $\hat{\bm{S}}_{\text{test}}$ is its estimate, and $L_{\text{test}}$ is the length of non-pilot test data units. As data-driven machine learning methods, all parameters (e.g., uncertainty quantification coefficients $\epsilon$’s) of combiners can be tuned using the popular cross-validation (e.g., one-shot cross-validation) method. The parameters can also be empirically tuned to save training times because cross-validation imposes a significant computational burden. This article mainly uses the empirical tuning method (i.e., trial-and-error) to tune each combiner to achieve its best average performance. For each test case, the MSE performances are averaged on $250$ Monte–Carlo episodes.
@@
    Since the signal model under impulse channel noises is no longer linear Gaussian, the optimal combiner in the MSE sense must be nonlinear. Therefore, the Kernel and the Kernel-DL methods have the potential to outperform other linear beamformers, i.e., to suppress outliers. However, in practice, the non-robust Kernel method may undergo numerical instability in calculating the inverse of the kernel matrix $\bm{K}$. Therefore, its actual MSEs are not necessarily smaller than those of linear beamformers. Nevertheless, the robust Kernel-DL method consistently outperforms all other beamformers.
@@
    Distributionally robust combiners (including diagonal-loading ones) can combat the adverse effect introduced by the limited pilot size and several types of uncertainties in the signal model (e.g., outliers). To be specific, for example, all diagonal-loading combiners can outperform their original non-diagonal-loading counterparts; cf. the Wiener and the Wiener-DL methods, the Wiener-CE and the Wiener-CE-DL methods, the Capon and the Capon-DL methods, and the Kernel and the Kernel-DL methods. In addition, the Wiener-DR beamformer (53) using the $F$-norm uncertainty set has the potential to outperform the Wiener-DL beamformer (30) that employs the simple uncertainty set (26).
@@
Remarks on Parameter Tuning: From experiments, we find that the uncertainty quantification coefficients $\epsilon$’s (e.g., in diagonal loading) can be neither too large nor too small. When $\epsilon$’s are too large, the combiners become overly conservative, while when $\epsilon$’s are too small, the combiners cannot offer sufficient robustness against data scarcity and model uncertainties. In both cases of inappropriate $\epsilon$’s, the performances of combiners degrade significantly. Therefore, $\epsilon$’s must be carefully tuned in practice, and a rigorous method to tune $\epsilon$’s can be the cross-validation method on the training data set (i.e., the pilot data set). If practitioners just pursue satisfaction rather than optimality, the empirical tuning method is recommended to save training times.
@@
Combiner MSE Time Combiner MSE Time
@@
Combiner MSE Time Combiner MSE Time
@@
Combiner MSE Time Combiner MSE Time
@@
Combiner MSE Time Combiner MSE Time
@@
Combiner MSE Time Combiner MSE Time
@@
Combiner MSE Time Combiner MSE Time